OpenStudy (anonymous):

A bag contains 12 red marbles, 5 yellow marbles, and 15 gree marbles. How many red marbles must be added to the 32 marbles already in the bag so that the probability of randomly selecting a red marble is 3/5?

jimthompson5910 (jim_thompson5910):

x = number of red marbles to add 12 red marbles already, so we will have 12+x red after we add x more red marbles 32 marbles total ----> 32+x total

jimthompson5910 (jim_thompson5910):

\[\Large \frac{12+x}{32+x} = \frac{3}{5}\] solve for x

OpenStudy (anonymous):

So the answer would be 18, right?

jimthompson5910 (jim_thompson5910):

yes it is

OpenStudy (anonymous):

Thank you!

jimthompson5910 (jim_thompson5910):

yw